# Diversity, equity, and inclusion for problems in additive number theory

### 19 Mar 2026 v2

@@ Abstract.
This is a survey of the diversity of problems in additive number theory. Equity requires the consideration of less currently popular problems, and suggests their inclusion in the additive canon. Of particular interest are problems about the sizes of sumsets of finite sets of integers and problems about the arithmetical structure of intersections of sumsets.
@@
Supported in part by PSC-CUNY Research Award Program grant 66197-00 54.
@@ 1. Canonical additive problems
Here are three of the most important (and most popular) areas of research in additive number theory:
@@ 2. Sum-product: A problem no one cared about
An important result, but did anyone pay attention? In 1991, Erdös wrote that “our paper with Szemerédi has nearly been forgotten.”
@@ 3. Freiman’s theorem
Almost no one had paid any attention to Grigorii Freiman’s remarkable work on inverse problems in additive number theory, even though he had published a book in Russian in 1966 and the American Mathematical Society had published an English translation in 1973. (Of course, in Freiman’s book and papers, the proof of his big theorem was hard to understand.)
Much later, various Hungarians worked on the theorem, and Ruzsa found a beautiful proof in 1994, which was exposed in the Springer Graduate Text Additive Number Theory: Inverse Theorems and the Geometry of Sumsets (Nathanson [13]).
Tim Gowers used this to obtain explicit estimates for Szemerédi’s theorem on $k$-term arithmetic progressions in dense sets of integers, and Freiman’s theorem and its consequences became another big business in additive number theory.
@@ 4. A noncanonical problem: Sizes of sumsets
In their 1983 paper on the sum-product conjecture, Erdős and Szemerédi wrote:
Let $2k-1\leq t\leq\frac{k^{2}+k}{2}$. It is easy to see that one can find a sequence of integers $a_{1}<\ldots<a_{k}$ so that there should be exactly $t$ distinct integers in the sequence $a_{i}+a_{j},1\leq i\leq j\leq k$.’
@@
The upper bound is combinatorial: If $A=\{a_{1},\ldots,a_{k}\}$ is a set of size $k$ whose $\binom{k}{2}$ sums $a_{i}+a_{j}$ with $a_{i}\neq a_{j}$ and $k$ sums $a_{i}+a_{i}$ are distinct, then the number of 2-fold sums is
@@
For example: The set $\{1,3,3^{2},\ldots,3^{k-1}\}$ has all 2-fold sums distinct.
For every set $A$ of integers, let
@@
For all $t\in\mathcal{R}(2,k)$, compute (construct) a set $A$ such that $|A|=k$ and $\left|2A\right|=t$.
Is there an algorithm?
@@
For all $t\in\mathcal{R}(2,k)$, there exists a set $A\subseteq\left[0,2^{k}-1\right]$ such that $|A|=k$ and $\left|2A\right|=t$, and $A$ can be computed in exponential time.
@@ 6. Missing sumset sizes for 3-fold sums of sets with 3 elements
Rajagopal proved the following fundamental result.
@@
For all $h$ there exists $k_{h}$ such that, for all $k>k_{h}$,
@@ 11. The moral of the story
It suggests one of Aesop’s fables. Aesop lived in ancient Greece between 620 and 564 BCE. Herodotus, in the fifth century BCE, mentioned him in his Histories in Book II, Chapter 134, where he referred to him as a fable writer and says he was a slave…. Aesop’s fables were transmitted orally and not collected and written down until three centuries after Aesop’s death. Stories have been added continuously and are still being added, so it is not clear which of Aesop’s fables were recited by Aesop.
One of the oldest and most famous of Aesop’s fables is variously known as “The Old Man and His Sons” or “The Bundle of Sticks”. It is numbered 53 in the Perry Index.
An old man had many sons who were always quarreling with one another. He had often begged them to live in harmony, but without success. One day he called them together, and producing a bundle of sticks, he asked them each to try and break it. Each tried with all his strength, but the bundle resisted their efforts. Then, cutting the cord that bound the sticks together, the old man told his sons to break them separately. This was done with the greatest ease. “See, my sons,” he exclaimed, “the power of unity! Bound together by brotherly love, you can defy almost every mortal ill; divided, you will fall a prey to your enemies.”

### 06 Apr 2026 v3

@@ Abstract.
There is a diversity of problems in additive number theory. Equity requires the consideration of less well known problems, and suggests their inclusion in the additive canon. Of particular interest are problems about the sizes of sumsets of finite sets of integers and problems about the arithmetical structure of intersections of sumsets.
@@
This is the text of a talk delivered at the New York Number Theory Seminar at the CUNY Graduate Center on February 19, 2026, and at the International Number Theory Conference (Alladi70) at the University of Florida on March 18, 2026.
@@ 1. Canonical additive problems
Here are three of the most important and best known research topics in additive number theory:
@@ 2. Sum-product: A problem no one cared about
An important result, but did anyone pay attention? In 1991, Erdös [5] wrote that “our paper with Szemerédi has nearly been forgotten.”
@@ 3. Freiman’s theorem
Almost no one had paid any attention to Grigorii Freiman’s remarkable work on inverse problems in additive number theory, even though he had published a book in Russian in 1966 and the American Mathematical Society (Freiman [9]) had published an English translation in 1973. (Of course, in Freiman’s book and papers, the proof of his big theorem was hard to understand.)
Much later, various Hungarians worked on the theorem, and Ruzsa [28] found a beautiful proof in 1994, which was exposed in the Springer Graduate Text Additive Number Theory: Inverse Theorems and the Geometry of Sumsets (Nathanson [16]).
Tim Gowers [10] used this to obtain explicit estimates for Szemerédi’s theorem on $k$-term arithmetic progressions in dense sets of integers, and Freiman’s theorem and its consequences became another big business in additive number theory.
@@ 4. A noncanonical problem: Sizes of sumsets
In their 1983 paper on the sum-product conjecture, Erdős and Szemerédi [6] wrote:
Let $2k-1\leq t\leq\frac{k^{2}+k}{2}$. It is easy to see that one can find a sequence of integers $a_{1}<\ldots<a_{k}$ so that there should be exactly $t$ distinct integers in the sequence $a_{i}+a_{j},1\leq i\leq j\leq k$.
@@
The upper bound is combinatorial: If $A=\{a_{1},\ldots,a_{k}\}$ is a set of size $k$ whose $\binom{k}{2}$ sums $a_{i}+a_{j}$ with $a_{i}\neq a_{j}$ and $k$ sums $a_{i}+a_{i}=2a_{i}$ are distinct, then the number of 2-fold sums is
@@
Such sets are called Sidon sets. For example, the set $\{1,3,3^{2},\ldots,3^{k-1}\}$ is a Sidon set.
For every set $A$ of integers, consider the 2-fold sumset
@@
For all $t\in\mathcal{R}(2,k)$, compute (construct) a set $A$ such that $|A|=k$ and $\left|2A\right|=t$. Is there an algorithm?
@@
For all $t\in\mathcal{R}(2,k)$, there exists a set $A\subseteq\left[0,2^{k}-1\right]$ such that $|A|=k$ and $\left|2A\right|=t$. The set $A$ can be computed in exponential time.
@@ 6. Missing sumset sizes for 3-fold sums of sets with 3 elements
The following result is fundamental.
@@
For all $h$, there exists $k_{h}$ such that, for all $k>k_{h}$,
@@ 11. The moral of the story
It suggests one of Aesop’s fables [1]. Aesop lived in ancient Greece between 620 and 564 BCE. Herodotus, in the fifth century BCE, mentioned him in his Histories in Book II, Chapter 134, where he referred to him as “Aesop the maker of fables” and writes that he was a slave of Iadmon, a Samian. Aesop’s fables were transmitted orally and not collected and written down until three centuries after Aesop’s death. Stories have been added continuously and are still being added, so it is not clear which of Aesop’s fables were recited by Aesop.
One of the oldest and most famous of Aesop’s fables is variously known as “The Old Man and His Sons” or “The Bundle of Sticks”. It is numbered 53 in the Perry Index [26].
There was an old man who had many sons. The sons were always fighting with each other, and the father kept trying unsuccessfully to get them to stop. One day he collected a large number of sticks, and, with a rope, tied them together into a thick bundle. He asked his sons to break the bundle. One by one they tried and failed. Then the father cut the rope and gave his sons the individual sticks, which they broke with ease. “Working together, you can succeed,” the father told them.